Weston Inc. had sales of $150,000 of sales, $75,000 of operating costs, and $10,000 of depreciation. The company borrowed long term $17,000 at a 7.0% interest rate, and its total tax rate was 35%. How much was the firm's net income?

$35,167.33

$37,018.24

$38,966.57

$41,476.50

$43,068.31
Answer #1

Interest =Loan*Interest Rate =17000*7% =1190
Net Income =(Sales-Operating Costs-Depreciation-Interest)*(1-Tax Rate)=(150000-75000-10000-1190)*(1-35%) =41476.50

Option d is correct option